how do you solve the problem x^2+y^2+12x-10y+45=0
using conic form circles

Answer by lwsshak3(11628)

how do you solve the problem x^2+y^2+12x-10y+45=0
using conic form circles
..
x^2+y^2+12x-10y+45=0
(x^2+12x+36)+(y^2-10y+25)=-45+36+25=16
(x+6)^2+(y-5)^2=16
This is an equation of a circle with center at (-6,5) with radius of 4